# Early Directional Convergence in Deep Homogeneous Neural Networks for Small Initializations

### v1

@@ 5.1 Early Directional Convergence
$\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}\geq\delta\eta/2,\text{ and }\frac{\mathbf{w}(\overline{T}_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}}\geq 1-\left(1+\frac{3}{\eta}\right)\epsilon,$
where ${\mathbf{u}}^{*}$ is a non-negative KKT  [^4] A KKT point $\hat{\mathbf{u}}$ of $\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{u})$ is referred to as non-negative KKT point if $\mathcal{N}_{\mathbf{z},\mathcal{H}}(\hat{\mathbf{u}})$ is non-negative. point of
$\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}_{\mathbf{y},\mathcal{H}}(\mathbf{u})=\mathbf{y}^{\top}\mathcal{H}(\mathbf{X};\mathbf{u}),$ (3)
or
$\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}\leq 2\delta\epsilon.$
@@
The above theorem first shows that for a given $\epsilon$, we can choose $\delta$ small enough such that for $t\in[0,T_{\epsilon}/\delta^{L-2}]$ the norm of the weights increases at most by a multiplicative factor. Thus, for small $\delta$ the weights will remain small for $t\in[0,T_{\epsilon}/\delta^{L-2}]$. Furthermore, since $L>2$, for smaller $\delta$ the weights will stay small for a longer duration of time. Next, the theorem describes what happens at $\overline{T}_{\epsilon}=T_{\epsilon}/\delta^{L-2}$, the duration for which the norm of the weights remain small. There are two possible scenarios. In the first scenario, the weights approximately converge in direction along a non-negative KKT point of (3), which is the constrained NCF defined with respect to $\mathbf{y}$ and neural network $\mathcal{H}$. Thus, in the early stages of training, the weights of the neural network converge in direction while their norm remains small. In the second scenario, the weights approximately converge to $\mathbf{0}$, which can be observed by noting that $\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}\leq 2\delta\epsilon$, where we can choose $\epsilon$ to be arbitrarily small. This is in contrast with the first scenario, where $\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}\geq\delta\eta/2$ and $\eta$ is a constant that does not depend on $\delta$. This shows that in the second scenario, compared to the first scenario, the weights become much smaller. The second scenario arises due to the convergence of the gradient dynamics of the NCF towards $\mathbf{0}$; see the proof for more details. We also note that for a fixed training data and neural network, $\mathbf{w}_{0}$ determines which of the above two scenarios will occur. In addition to this, the positive constant $\eta$ in the above theorem is also determined by $\mathbf{w}_{0}$ and is independent of $\delta$.
5.1.1 Proof Sketch of Theorem 5.1
@@
We next briefly provide proof sketch of the above theorem. We begin by describing the (positive) gradient flow dynamics of the NCF. Recall, for a given vector $\mathbf{z}$ and a neural network $\mathcal{H}$, the NCF is defined as
The gradient flow resulting from maximizing the above function is
$\frac{d{\mathbf{u}}}{dt}=\nabla\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{u}),{\mathbf{u}}(0)=\mathbf{u}_{0},$ (5)
@@
Lemma 5.2.
Under 1, for any solution $\mathbf{u}(t)$ of (5), one of the following two possibilities is true.
@@
    There exists some finite $T^{*}$ such that $\lim_{t\rightarrow T^{*}}\|\mathbf{u}(t)\|_{2}=\infty$. Further, $\lim_{t\rightarrow T^{*}}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ exists and is equal to a non-negative KKT point of the optimization problem
@@
    For all $t\geq 0$, $\|\mathbf{u}(t)\|_{2}<\infty,$ and either $\lim_{t\rightarrow\infty}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ exists or $\lim_{t\rightarrow\infty}{\mathbf{u}(t)}=0.$ If $\lim_{t\rightarrow\infty}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ exists then it is equal to non-negative KKT point of (6).
To prove directional convergence in the above lemma, we require the neural networks to be definable with respect to an o-minimal structure. Such a requirement allows us to use the unbounded version of Kurdyka-Lojasiewicz inequality proved in [31] for establishing directional convergence. A similar result was demonstrated in [1] for two-homogeneous networks. However, in that instance, the solution $\mathbf{u}(t)$ remains finite for all finite time [^5] See Lemma D.1., whereas in our case, $\mathbf{u}(t)$ may become unbounded at some finite time. We next explain our proof technique for Theorem 5.1 assuming square loss is used for training. The evolution of $\mathbf{w}(t)$ is governed by the differential equation
@@
where (as mentioned) $\delta>0$, and recall $\mathcal{J}(\mathbf{X};\mathbf{w}):\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$ is the Jacobian of the function $\mathcal{H}(\mathbf{X};\mathbf{w})$ with respect to $\mathbf{w}$. We define $\mathbf{s}(t)=\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)$; then, $\mathbf{s}(0)=\mathbf{w}_{0}$. Further, by using the $L$-homogeneity of $\mathcal{H}\left(\mathbf{X};\mathbf{w}\right)$ and $(L-1)$-homogeneity of $\mathcal{J}(\mathbf{X};\mathbf{w})$ (Lemma B.1), we can derive the evolution of $\mathbf{s}(t)$ as follows
@@
$\frac{d\mathbf{s}}{dt}$ $=\mathcal{J}\left(\mathbf{X};\mathbf{s}\right)^{\top}(\mathbf{y}-\delta^{L}\mathcal{H}\left(\mathbf{X};\mathbf{s}\right)),\mathbf{s}(0)=\mathbf{w}_{0}.$ (8)
@@
Now, we compare the dynamics for $\mathbf{s}(t)$ in (8) with the dynamics of gradient flow of the NCF defined with respect to $\mathbf{y}$ and neural network $\mathcal{H}$,
$\dot{\mathbf{u}}=\nabla\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{u})=\mathcal{J}(\mathbf{X};\mathbf{u})^{\top}\mathbf{y},{\mathbf{u}}(0)=\mathbf{w}_{0}.$ (9)
@@
Observe that if $\delta=0$, then the two dynamics are exactly the same, and since $\mathbf{s}(0)=\mathbf{w}_{0}=\mathbf{u}(0)$, we have that $\mathbf{s}(t)$ and $\mathbf{u}(t)$ will also be same. For $\delta>0$, there is an extra term $\delta^{L}\mathcal{H}\left(\mathbf{X};\mathbf{s}\right)$ present in (8). For small $\delta$, we show that this extra term is small and while $\mathbf{s}(t)$ and $\mathbf{u}(t)$ are not exactly the same, the difference between them is small, i.e., $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}$ is small. Moreover, we show that the difference can be made sufficiently small for a sufficiently long time by choosing $\delta$ small enough. We combine this fact with Lemma 5.2 to prove Theorem 5.1. Specifically, from Lemma 5.2, we know that $\mathbf{u}(t)$ will either converge to $\mathbf{0}$ or converge in direction to a KKT point of the constrained NCF. Therefore, we first choose a time (say $T$) such that $\mathbf{u}(T)$ has approximately converged in direction to a KKT point of the constrained NCF or is in the neighborhood of $\mathbf{0}$. Then, we choose $\delta$ sufficiently small such that $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}$ is small for all $t\in[0,T]$. Since $\mathbf{s}(t)=\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)$, we have that $\|\frac{1}{\delta}\mathbf{w}\left(\frac{T}{\delta^{L-2}}\right)-\mathbf{u}(T)\|_{2}$ is also small. Finally, since $\delta$ is a positive scalar, and $\mathbf{u}(T)$ approximately converges in direction to a KKT point of the constrained NCF or is approximately $\mathbf{0}$, it follows that $\mathbf{w}\left(\frac{T}{\delta^{L-2}}\right)$ also approximately converges in direction to a KKT point of the constrained NCF or is approximately $\mathbf{0}$. As a quick aside, we note that there are two key aspects in the definition $\mathbf{s}(t)=\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)$. First, we have divided by $\delta$. We know that $\delta$ controls the scale of initialization and is assumed to be small. Dividing by $\delta$ removes that scaling and makes $\mathbf{s}(t)$ a constant scale vector. More importantly, dividing by $\delta$ does not alter the direction, i.e., the direction of $\mathbf{s}(t)$ and $\mathbf{w}(t)$ are always the same. Thus, establishing directional convergence of $\mathbf{s}(t)$ will also imply directional convergence of $\mathbf{w}(t)$. Second, we rescale the time by $\delta^{L-2}$. To understand the importance of this time rescaling, let us look at the magnitude of the gradient of the loss at initialization in terms of $\delta$. Since $\delta$ is small and $\mathbf{w}_{0}$ has unit norm, $\|\mathcal{J}(\mathbf{X};\mathbf{w}(0))^{\top}\left(\mathbf{y}-\mathcal{H}(\mathbf{X};\mathbf{w}(0))\right)\|_{2}\approx\|\mathcal{J}(\mathbf{X};\mathbf{w}(0))^{\top}\mathbf{y}\|_{2}=O(\delta^{L-1})$, where in the last equality we used $(L-1)$-homogeneity of the Jacobian. We observe that for initialization of scale $\delta$, the gradient scales as $\delta^{L-1}$. Thus, in the early stages the gradient will have negligible effect on the weights, in terms of both magnitude and direction. Moreover, based on the relative scaling of the gradient and the initialization, the gradient will only start impacting the weights after $O(1/\delta^{L-2})$ time has elapsed. Hence, for smaller initialization, the gradient flow will take longer to make an impact on the weights. Thus, in a way, rescaling time by $\delta^{L-2}$ brings the time of $\mathbf{s}(t)$ to a “constant scale”. We also highlight the importance of homogeneity. It allows the above two changes to interact nicely with the gradient flow and results in dynamics of $\mathbf{s}(t)$ that are essentially independent of $\delta$, provided $\delta$ is small. Finally, in comparison to [1], which studies early directional convergence for two-homogeneous neural networks, the main conceptual innovation in this paper is the time rescaling by $\delta^{L-2}$. In fact, the proof in that work relies on showing that the difference between $\mathbf{w}(t)/\delta$ and $\mathbf{u}(t)$ remains small for a sufficiently long time. To see a motivation for this, note that we can get (8) from (7) for $L=2$ as well. However, for $L=2$, the time rescaling factor $\delta^{L-2}$ will become unity, yielding $\mathbf{s}(t)=\mathbf{w}(t)/\delta$. Now, since the gradient flow dynamics for the NCF of two-homogeneous neural networks also converges to a KKT point of the constrained NCF, to establish early directional convergence one must show that the difference between $\mathbf{w}(t)/\delta$ and $\mathbf{u}(t)$ remains small for a sufficiently long time. Another notable difference with [1] is in the gradient flow dynamics of the NCF. As mentioned previously, from Lemma 5.2, we know that for $L>2$, the gradient flow dynamics of the NCF can become unbounded at some finite time, whereas for $L=2$, the dynamics remains finite for finite time. This difference necessitates a more careful analysis of the early stages of training in order to ensure directional convergence for deep homogeneous neural networks.

### v2

@@ 4.1 Main Result
$\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}\geq\delta\eta/2,\text{ and }\mathbf{u}_{*}^{\top}\mathbf{w}(\overline{T}_{\epsilon})/{\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}}\geq 1-\left(1+{3}/{\eta}\right)\epsilon,$
where $\mathbf{u}_{*}$ is a non-negative KKT point of
$\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}_{-\ell^{\prime}(\mathbf{0},\mathbf{y}),\mathcal{H}}(\mathbf{u})=\max_{\|\mathbf{u}\|_{2}^{2}=1}-\ell^{\prime}(\mathbf{0},\mathbf{y})^{\top}\mathcal{H}(\mathbf{X};\mathbf{u}),$ (3)
or $\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}\leq 2\delta\epsilon.$
The above theorem first shows that for a given $\epsilon$, we can choose $\delta$ small enough such that for $t\in[0,T_{\epsilon}/\delta^{L-2}]$ the norm of the weights increases at most by a multiplicative factor. Thus, for small $\delta$ the weights will remain small for $t\in[0,T_{\epsilon}/\delta^{L-2}]$. Furthermore, since $L>2$, for smaller $\delta$ the weights will stay small for a longer duration of time.
@@
Next, the theorem describes what happens at $\overline{T}_{\epsilon}=T_{\epsilon}/\delta^{L-2}$, the duration for which the norm of the weights remain small. There are two possible scenarios. In the first scenario, the weights approximately converge in direction along a non-negative KKT point of eq. 3, which is the constrained NCF defined with respect to $-\ell^{\prime}(\mathbf{0},\mathbf{y})$ and neural network $\mathcal{H}$. Thus, in the early stages of training, the weights of the neural network converge in direction while their norm remains small. In the second scenario, the weights approximately converge to $\mathbf{0}$, which can be observed by noting that $\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}\leq 2\delta\epsilon$, where we can choose $\epsilon$ to be arbitrarily small. This is in contrast with the first scenario, where $\|\mathbf{w}(\overline{T}_{\epsilon})\|_{2}\geq\delta\eta/2$ and $\eta$ is a constant that does not depend on $\delta$. This shows that in the second scenario, compared to the first scenario, the weights become much smaller. The second scenario arises due to the convergence of the gradient dynamics of the NCF towards $\mathbf{0}$; see the proof for more details. We also note that for a fixed training data and neural network, $\mathbf{w}_{0}$ determines which of the above two scenarios will occur. In addition to this, the positive constant $\eta$ in the above theorem is also determined by $\mathbf{w}_{0}$ and is independent of $\delta$.
It can be shown that for square loss, $\ell(\hat{y},y)=\frac{1}{2}(\hat{y}-y)^{2}$ and $-\ell^{\prime}(0,\mathbf{y})=\mathbf{y}$, and for logistic loss, $\ell(\hat{y},y)=\ln(1+e^{-\hat{y}y})$ and $-\ell^{\prime}(0,\mathbf{y})=\mathbf{y}/2$. Therefore, for square and logistic loss, the objective function of the constrained NCF in eq. 3 will be $\mathbf{y}^{\top}\mathcal{H}(\mathbf{X};\mathbf{u})$. We also study KKT points of the constrained NCF for feed-forward neural networks in Section 5, which some readers may find helpful.
Overall, the phenomenon of early directional convergence in deep homogeneous networks is similar to two-homogeneous networks (Kumar & Haupt, 2024, Theorem 5.1), except for one difference. The amount of time required for directional convergence in deep homogeneous networks has an additional factor of $1/\delta^{L-2}$. Thus, directional convergence takes longer as depth increases.
4.1.1 Proof Sketch of Theorem 1
@@
We next provide a brief proof sketch of the above theorem. We begin by describing the (positive) gradient flow dynamics of the NCF.
For a given vector $\mathbf{z}$ and a neural network $\mathcal{H}$, the gradient flow resulting from maximizing the NCF is
$\dot{{\mathbf{u}}}=\nabla\mathcal{N}_{\mathbf{z},\mathcal{H}}(\mathbf{u}),{\mathbf{u}}(0)=\mathbf{u}_{0},$ (4)
@@
Lemma 2.
Under Assumption 1, for any solution $\mathbf{u}(t)$ of eq. 4, one of the following two possibilities is true.
@@
    There exists some finite $T^{*}$ such that $\lim_{t\rightarrow T^{*}}\|\mathbf{u}(t)\|_{2}=\infty$. Further, $\lim_{t\rightarrow T^{*}}{\mathbf{u}(t)}/{\|\mathbf{u}(t)\|_{2}}$ exists and is equal to a non-negative KKT point of the optimization problem
@@
    For all $t\geq 0$, $\|\mathbf{u}(t)\|_{2}<\infty,$ and either $\lim_{t\rightarrow\infty}{\mathbf{u}(t)}/{\|\mathbf{u}(t)\|_{2}}$ exists or $\lim_{t\rightarrow\infty}{\mathbf{u}(t)}=0.$ If $\lim_{t\rightarrow\infty}{\mathbf{u}(t)}/{\|\mathbf{u}(t)\|_{2}}$ exists then it is equal to non-negative KKT point of eq. 5.
To prove the above lemma, we require the neural networks to be definable with respect to an o-minimal structure. Such a requirement allows us to use the unbounded version of Kurdyka-Lojasiewicz inequality proved in Ji & Telgarsky (2020) for establishing directional convergence. A similar result was demonstrated in Kumar & Haupt (2024) for two-homogeneous networks. However, in that instance, the solution $\mathbf{u}(t)$ remains finite for all finite time (see Lemma 18), whereas in our case, $\mathbf{u}(t)$ may become unbounded at some finite time. Also, for our proof, we had to derive the rate at which $\mathbf{u}(t)$ grows, showing that, if $\mathbf{u}(t)$ becomes unbounded, then $\|\mathbf{u}(t)\|_{2}\geq\kappa/(T^{*}-t)^{1/(L-2)}$, for some $\kappa>0$ and for all $t$ near $T^{*}$.
We next explain our proof technique for Theorem 1 assuming square loss is used for training. The evolution of $\mathbf{w}(t)$ is governed by the differential equation
@@
where (as mentioned) $\delta>0$, and recall $\mathcal{J}(\mathbf{X};\mathbf{w}):\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$ is the Jacobian of the function $\mathcal{H}(\mathbf{X};\mathbf{w})$ with respect to $\mathbf{w}$. We define $\mathbf{s}(t)=\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)$; then, $\mathbf{s}(0)=\mathbf{w}_{0}$. Further, by using the $L$-homogeneity of $\mathcal{H}\left(\mathbf{X};\mathbf{w}\right)$ and $(L-1)$-homogeneity of $\mathcal{J}(\mathbf{X};\mathbf{w})$ (Lemma 9), we can derive the evolution of $\mathbf{s}(t)$ as follows
@@
$\dot{\mathbf{s}}=\mathcal{J}\left(\mathbf{X};\mathbf{s}\right)^{\top}(\mathbf{y}-\delta^{L}\mathcal{H}\left(\mathbf{X};\mathbf{s}\right)),\mathbf{s}(0)=\mathbf{w}_{0}.$ (7)
@@
Now, we compare the dynamics for $\mathbf{s}(t)$ in eq. 7 with the dynamics of gradient flow of the NCF defined with respect to $\mathbf{y}$ and neural network $\mathcal{H}$,
$\dot{\mathbf{u}}=\nabla\mathcal{N}_{\mathbf{y},\mathcal{H}}(\mathbf{u})=\mathcal{J}(\mathbf{X};\mathbf{u})^{\top}\mathbf{y},{\mathbf{u}}(0)=\mathbf{w}_{0}.$ (8)
Observe that if $\delta=0$, then the two dynamics are exactly the same, and since $\mathbf{s}(0)=\mathbf{w}_{0}=\mathbf{u}(0)$, we have that $\mathbf{s}(t)$ and $\mathbf{u}(t)$ will also be same. For $\delta>0$, there is an extra term $\delta^{L}\mathcal{H}\left(\mathbf{X};\mathbf{s}\right)$ present in eq. 7. For small $\delta$, we show that this extra term is small and while $\mathbf{s}(t)$ and $\mathbf{u}(t)$ are not exactly the same, the difference between them is small, i.e., $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}$ is small. Moreover, we show that the difference can be made sufficiently small for a sufficiently long time by choosing $\delta$ small enough. We combine this fact with Lemma 2 to prove Theorem 1. Specifically, from Lemma 2, we know that $\mathbf{u}(t)$ will either converge to $\mathbf{0}$ or converge in direction to a KKT point of the constrained NCF. Therefore, we first choose a time (say $T$) such that $\mathbf{u}(T)$ has approximately converged in direction to a KKT point of the constrained NCF or is in the neighborhood of $\mathbf{0}$. Then, we choose $\delta$ sufficiently small such that $\|\mathbf{s}(t)-\mathbf{u}(t)\|_{2}$ is small for all $t\in[0,T]$. Since $\mathbf{s}(t)=\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)$, we have that $\|\frac{1}{\delta}\mathbf{w}\left(\frac{T}{\delta^{L-2}}\right)-\mathbf{u}(T)\|_{2}$ is also small. Finally, since $\delta$ is a positive scalar, and $\mathbf{u}(T)$ approximately converges in direction to a KKT point of the constrained NCF or is approximately $\mathbf{0}$, it follows that $\mathbf{w}\left(\frac{T}{\delta^{L-2}}\right)$ also approximately converges in direction to a KKT point of the constrained NCF or is approximately $\mathbf{0}$.
@@
As a quick aside, we note that there are two key aspects in the definition $\mathbf{s}(t)=\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)$. First, we have divided by $\delta$. We know that $\delta$ controls the scale of initialization and is assumed to be small. Dividing by $\delta$ removes that scaling and makes $\mathbf{s}(t)$ a constant scale vector. More importantly, dividing by $\delta$ does not alter the direction, i.e., the direction of $\mathbf{s}(t)$ and $\mathbf{w}(t)$ are always the same. Thus, establishing directional convergence of $\mathbf{s}(t)$ will also imply directional convergence of $\mathbf{w}(t)$. Second, we rescale the time by $\delta^{L-2}$. To understand the importance of this time rescaling, let us look at the magnitude of the gradient of the loss at initialization in terms of $\delta$. Since $\delta$ is small and $\mathbf{w}_{0}$ has unit norm, $\|\mathcal{J}(\mathbf{X};\mathbf{w}(0))^{\top}\left(\mathbf{y}-\mathcal{H}(\mathbf{X};\mathbf{w}(0))\right)\|_{2}\approx\|\mathcal{J}(\mathbf{X};\mathbf{w}(0))^{\top}\mathbf{y}\|_{2}=O(\delta^{L-1})$, where in the last equality we used $(L-1)$-homogeneity of the Jacobian. We observe that for initialization of scale $\delta$, the gradient scales as $\delta^{L-1}$. Thus, in the early stages the gradient will have negligible effect on the weights, in terms of both magnitude and direction. Moreover, based on the relative scaling of the gradient and the initialization, the gradient will only start impacting the weights after $O(1/\delta^{L-2})$ time has elapsed. Hence, for smaller initialization, the gradient flow will take longer to make an impact on the weights. Thus, in a way, rescaling time by $\delta^{L-2}$ brings the time of $\mathbf{s}(t)$ to a “constant scale”. We also highlight the importance of homogeneity. It allows the above two changes to interact nicely with the gradient flow and results in dynamics of $\mathbf{s}(t)$ that are essentially independent of $\delta$, provided $\delta$ is small.
Finally, in comparison to Kumar & Haupt (2024), which studies early directional convergence for two-homogeneous neural networks, the main conceptual innovation in this paper is the time rescaling by $\delta^{L-2}$. In fact, the proof in that work relies on showing that the difference between $\mathbf{w}(t)/\delta$ and $\mathbf{u}(t)$ remains small for a sufficiently long time. To see a motivation for this, note that we can get eq. 7 from eq. 6 for $L=2$ as well. However, for $L=2$, the time rescaling factor $\delta^{L-2}$ will become unity, yielding $\mathbf{s}(t)=\mathbf{w}(t)/\delta$. Now, since the gradient flow dynamics for the NCF of two-homogeneous neural networks also converges to a KKT point of the constrained NCF, to establish early directional convergence one must show that the difference between $\mathbf{w}(t)/\delta$ and $\mathbf{u}(t)$ remains small for a sufficiently long time. Another notable difference with Kumar & Haupt (2024) is in the gradient flow dynamics of the NCF. As mentioned previously, from Lemma 2, we know that for $L>2$, the gradient flow dynamics of the NCF can become unbounded at some finite time, whereas for $L=2$, the dynamics remains finite for finite time. This difference necessitates a more careful analysis of the early stages of training in order to ensure directional convergence for deep homogeneous neural networks.
Challenges for non-differentiable neural networks:
Theorem 1 holds for neural networks with locally Lipschitz gradient, and thus excludes ReLU neural networks. We briefly describe the challenges in extending Theorem 1 to ReLU networks in Appendix D.